Ticket Number: HEA-NOC/20100616-3 Ticket Status: CLOSE
Ticket Type: unscheduled Resolver: LanComms
Ticket Opened: 20100616 11:45 UTC+1 Problem Start: 20100616 18:46 UTC+1
Ticket Closed: 20100618 09:01 UTC+1 Problem End: 20100617 16:13 UTC+1
Site/line: Blanch Services Network
Switch problem with blanch-mgt1. All traffic on ports Fa0/25 - Fa0/48 are
not transmitting packets/bits.
Traffic to PDUs connected to this management switch are unreachable at the
moment: 302, 30A, 308, 307, 306, 305, 304, 30F, 30E, 30D, 30C and 30B,
despite the interfaces showing as "up up".
20100616 11:33 OMG The trunk link to blanch-sr2 was not blocking traffic
on the trunk link from blanch-mgt1 so was shut down. JB has logged a call
20100616 14:17 JB - Lancomms are to replace the switch and expect to be
on-site with the new switch within 90 mins.
20100616 15:39 JB - Lancomms are on-site and will be starting the switch
swap in 10 minutes.
20100616 16:17 JB - Switch swap complete. All interfaces are now passing
traffic again and the STP problems have disappeared. Monitoring for 24h
20100618 09:00 JB - No issues with the new blanch-mgt1 switch. Closing
Time to Fix:
21 hours, 30 mins
This ticket can be monitored at http://www.hea.net/tickets/20100616-3
HEAnet Limited. Registered in Ireland, No. 275301.